OpenStudy (shadowcain):

For which operations is the set {–1, 0, 1} closed? Choose all answers that are correct. A. addition B. division C. multiplication D. subtraction

OpenStudy (shadowcain):

Division and Multiplication, right?

OpenStudy (shadowcain):

@zzr0ck3r @mathmate @King.Void. @SolomonZelman

OpenStudy (blackstreet23):

what does it mean set closed?

OpenStudy (anonymous):

A set of elements is closed under an operation if, when you apply the operation to elements of the set, you always get another element of the set.

OpenStudy (anonymous):

It's either A and D or all of them. I'm not sure

OpenStudy (zzr0ck3r):

ok let us go through them A) Is there a way to add two elements from the set and not get an element in the set?

OpenStudy (shadowcain):

Oh wait it can't be division because you can't divide by 0.....

OpenStudy (zzr0ck3r):

not addition because 1+1 = 2 is not in the set

OpenStudy (shadowcain):

yes -1 + 1 = 0

OpenStudy (shadowcain):

right?

jimthompson5910 (jim_thompson5910):

while `-1 + 1 = 0` is true, the fact that `1+1=2` means that you can take the elements `1` and `1` (repeat selections are possible), add them up, and get a number that is NOT in the set. So this set is NOT closed under addition

OpenStudy (shadowcain):

0 x -1 = 0 0 x 1 = 0 1 - 1 = 0 So, is it multiplication and subtraction?

jimthompson5910 (jim_thompson5910):

make a 3x3 table 3 rows for the elements -1,0,1 3 columns for the elements -1,0,1 

jimthompson5910 (jim_thompson5910):

Now let's say we will use the subtraction operation to fill out the table, we subtract the value on the left from the value up top eg: -1 minus -1 = -1+1 = 0 and this goes in the first row, first column the rest of the table is filled out the same way